OpenStudy (anonymous):

What is 6x^3/4 in radical form?

jimthompson5910 (jim_thompson5910):

\[\LARGE 6x^{\frac{3}{4}}\] \[\LARGE 6\sqrt[4]{x^3}\] So \[\LARGE 6x^{\frac{3}{4}} = 6\sqrt[4]{x^3}\]